# Conformer-Based Speech Recognition On Extreme Edge-Computing Devices

### v1

@@ 1 Introduction
Conformer-based (Gulati et al. 2020) end-to-end (E2E) automatic speech recognition (ASR) (Yao et al. 2021; Zhang et al. 2022) with streaming capabilities (He et al. 2019) have made numerous advances recently. This has paved the way for fully neural speech recognition on resource-constrained mobile devices. These systems also have numerous advantages over conventional hybrid hidden Markov model-based ASR (Hinton et al. 2012).
First, the training procedure is simplified; the entire system can be defined in a single deep learning framework such as PyTorch or TensorFlow. Second, recent work (e.g. Miao et al. 2019; Sainath et al. 2020; Li et al. 2020) shows E2E ASR systems can provide better Word-Error-Rate (WER) when compared to conventional hybrid ASR systems. Third, with the continued advancement of deep learning applications, special hardware accelerators such as NVIDIA’s Graphics Processing Units (GPU), Google’s Tensor Processing Units (TPU), and Apple’s Neural Engine (ANE) are becoming increasingly popular. A fully neural ASR system can best utilize such hardware advancements and operate with high throughput while minimizing energy consumption. Conventional hybrid ASR systems typically require a computationally-expensive decoder to properly search the language model. To improve decoder efficiency, ASR on mobile devices typically uses static decoders, like those built on Weighted Finite State Transducers (WFSTs) (Mohri, Pereira, and Riley 2002). WFST-based decoders are more efficient than fully-dynamic decoders, but remain a computational bottleneck in conventional mobile device large vocabulary continuous speech recognition.
In this paper, we present optimizations to enable E2E fully neural network based ASR system under resource-constrained environments, such as smartphones, wearables, and home automation devices. Operating fully offline saves cloud computing resources while providing stronger user privacy guarantees, as the user’s speech does not need to be transmitted outside of the device.
@@ 2 Prior Work
Improving the efficiency of the Transformer architecture has seen substantial interest. Tay et al. (2023) provides a comprehensive survey primarily concentrating on improvements to the model architecture. Kim et al. (2023) is another noteworthy resource , which delves deeper into considerations specific to hardware configurations. Linear Transformer (Katharopoulos et al. 2020) is a key technique, mitigating the computationally expensive softmax function (Bridle 1989) within the attention mechanism. Softmax is also susceptible to numeric overflow problems when computing with limited numerical range. Hoffer et al. (2018); Zhang and Sennrich (2019) discuss alternative normalization methods other than Batchnorm (Ioffe and Szegedy 2015) and Layernorm (Ba, Kiros, and Hinton 2016) to improve computational efficiency and numerical stability in low precision environments. Principles for optimizing transformers have been described in Apple (2022) which target Apple hardware, but are generally applicable for similar devices. Within the domain of speech recognition, Squeezeformer (Kim et al. 2022) stands as a seminal work focusing on efficiency optimization, particularly with respect to the Conformer architecture. The paper uses depthwise separable convolution subsampling to substantially save computation which is central to MobileNet (Howard et al. 2017). It’s worth mentioning that the majority of prior work focuses on improving training efficiency by making modifications to the existing model architecture. As a result, these changes require model retraining to achieve efficiency improvements. In contrast, our research primarily concentrates on post-training, inference-only processes while avoiding model retraining whenever possible.
3 Model Architecture
Our foundational model architecture is built upon the Conformer connectionist temporal classification (Graves et al. 2006) (CTC) ASR system. We’ve additionally incorporated an architectural enhancement by substituting the vanilla 2D convolution with depthwise separable convolution. This change reduces the model’s computational needs, and particularly fits well with the inference capabilities of mobile devices.
3.1 Conformer CTC
Acoustic Encoder
@@ 3.1 Conformer CTC
Similar to prior work (e.g. Gulati et al. 2020), we stack transformer (Vaswani et al. 2017) layers and convolution (LeCun et al. 1998) layers alternatively to convert speech frames into high-level representation. Transformer is responsible for capturing long-term dependency while convolution takes care of local patterns. We use a relative sinusoidal positional encoding (Dai et al. 2019) into the transformer layers. Since our goal is to stream ASR on edge devices, we adopt the chunk-based attention strategy to better balance accuracy and dependency of future audio frames (Vaswani et al. 2017). The key concept is to group the adjacent frames into a chunk and allow attention to access history frames as well as frames in the same chunk.
Objective Functions
Connectionist Temporal Classification (CTC) (Graves et al. 2006), Attention-based Encoder Decoder (AED) (Chan et al. 2016), and Transducer (Graves 2012) are the three mainstream objective functions and recognition algorithms for fully neural ASR. We follow Yao et al. (2021); Zhang et al. (2022) to train our model with the multitask objective of CTC and AED. We use only CTC for decoding. Because the optimization described next is for the acoustic encoder, it is applicable to other combinations of objective functions and recognition algorithms.
@@ 3.2 Depthwise Separable Convolution
In the original Conformer encoder design (Gulati et al. 2020), the subsampling module at the beginning of the architecture is implemented using two vanilla convolution layers. Our profiling shows that vanilla convolution subsampling accounts for 32.8% of the overall computation and becomes expensive on resource-constrained devices. To alleviate this bottleneck, we used the idea of depthwise separable convolution (Howard et al. 2017; Chollet 2017) as a drop-in replacement and reduced this computational bottleneck to 4.0% whilst maintaining the WER (Kim et al. 2022).
While most of the research emphasizes depthwise separable convolution’s (DWS) computational efficiency and small memory footprint, its effect on reducing dynamic range of the outputs needs more study. The possible reason could be that DWS reduces the number of multiply-accumulate operations needed for the convolution filters, hence the chance of bigger values. Low numeric range is of great importance for model deployment on edge devices equipped with hardware accelerators. Those hardware often operate in low precision (e.g.16bit floating point) to ease the burden of storage and memory and are exposed to overflow.
4 Performance Optimizations
In this section, we describe specific optimizations we implemented to achieve high performance execution of our modified Conformer model on smartphones and wearable devices. In Section 5, we will assume that our target accelerator is Apple Neural Engine (ANE). For other sections, such an assumption will not be held as the optimizations are applicable to hardware accelerators in general.
@@ 4.1 Memory-aware Graph Execution
In Apple (2022), four principles are elaborated for optimizing transformers on the Apple Neural Engine:
@@
The key idea behind these 4 principles is being aware of high cost invoked by memory copies between CPU and our hardware accelerator. In our implementation, we adhered to the aforementioned principles. Firstly, we transposed the input and output of Conformer CTC, expanding the input tensor to the desired shape of $(B,C,1,S)$. This transformation allowed us to execute most layers on the hardware accelerator as per Principle 1. Additionally, we extensively employed split and concatenation operations to enhance L2 cache residency (Principle 2). To address the issue of undesired memory copies resulting from batched matrix multiplication layers, we replaced them with Einstein summation operations (Principle 3). Finally, operations not supported by hardware accelerator were positioned at the beginning or end of the network graph, thus minimizing copies in the memory.
@@ 4.2 Stability of Layer Normalization
Layer normalization has become the de facto normalization method in transformers after Attention is all you need (Vaswani et al. 2017). This normalization technique is widely used in the Conformer CTC architecture. On the other hand, modern hardware accelerators for deep learning often exploit lower precision compute paths in order to reduce memory and boost computation throughput. In the Conformer model, we observed that layer normalization and hardware accelerators are often in dissonance with each other. The reason is that skip connections in the Conformer model join values of varying magnitudes to a single tensor and this often leads to numerical underflows or overflows in low precision compute paths. For example, the maximum value is 65504 in half precision floating point format (IEEE 2008). As a contrast, the maximum value is $383.4e$ 38 in single precision floating point format.
@@
Definition 1. A desirable low precision pre-normalizer maps a distribution of values to a bounded region, $[-C,C]$, for some small $C$.
Lemma 1. $f(\textbf{x})=\frac{\textbf{x}}{\frac{1}{n}\sum_{i=1}^{n}|x_{i}|}$ is a desirable low precision pre-normalizer for uniform distributions.
[…]
Thus, $f(\textbf{x})=\frac{2\textbf{x}}{L}\sim unif[-2,2]$.
Lemma 2. $f(\textbf{x})=\frac{\textbf{x}}{\frac{1}{n}\sum_{i=1}^{n}|x_{i}|}$ is a desirable low precision pre-normalizer for normal distributions.
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n}|x_{i}|$ $=\mathbb{E}[|\textbf{x}|]$
[…]
$=\sqrt{\frac{2}{\pi}}\sigma.$
@@
$=\Big{(}\frac{|-\frac{S}{2}|}{\frac{1}{2}(\frac{2}{M})^{1/p}\sum_{i=1}^{n}|v_{i}|}\Big{)}^{p}+\Big{(}\frac{|\frac{S}{2}|}{\frac{1}{2}(\frac{2}{M})^{1/p}\sum_{i=1}^{n}|v_{i}|}\Big{)}^{p}$ (3)
@@
On a practical note, the pre-normalizer we used for our experiment was the one from Lemmas 1 and 2 with $n=512$, which gave a slightly lower normalization constant than what Corollary 1 suggests. This worked well in our setup because attaining or even getting close to the maximum value as stated in Lemma 3 requires atypical distribution of values with very few extreme values and everything else being 0. This does not happen in practice, however, with the most common distribution of values observed being Gaussian.
@@ 4.3 Scaling of Softmax
Another common constraint on hardware accelerators is their limited support in complex operations. For example, hardware accelerators may choose to omit support for exponential operations (Hu et al. 2018; Li et al. 2018). In such cases, we seek to implement such operations in memory instead, namely using lookup tables (LUT). However, since LUTs are slow and expensive in terms of memory consumption, we would like the tables to be as small as possible. To this end, we introduce a technique called conditional re-scaling for softmax layers:
@@
To interpret the above transformation, we first assume that our LUT gives reasonably accurate approximation for $x_{i}$’s below 4096. Next we take FP16 as an example of our low precision compute paths. We note that for values greater than 4096, gaps between values jump in increments of 4 according to IEEE 754-2008 (IEEE 2008). Under such scenario, the softmax function behaves similarly to an argmax operation. Since gaps of values between 2048 and 4096 jump in increments of 2, the “argmax behavior” is largely preserved after the re-scaling and exponentiation.
@@
Figure 2: Energy consumption (in joules) for 200 queries of the original Conformer CTC vs Depthwise Separable Convolution (DWS) architectures. Blue and green bars represent the values on CPU and hardware accelerators, respectively. Note that the y-axis is in log scale.
@@ 5.1 Setup
The training corpus contains 17k-hour audio-transcript pairs where the audio is randomly sampled from anonymized virtual assistant queries and the corresponding transcripts are human-annotated. We curate 20k queries in the same manner to form an accuracy test set. We use it to examine the accuracy of the optimizations. 200 queries are sampled from the accuracy test set and serve as the performance test set. We want to rule out as many computationally intensive components as possible during speech recognition, so the audio is decoded with CTC prefix beam search (Graves et al. 2006). The data source and the training recipe do not play important role in the experiments because the proposed methods focus on hardware acceleration. The experiments are conducted on iPhone XR and Apple Watch Series 7.
Two models (conv2d6 and dws2d6) are trained with the same hyper-parameters but minor difference in subsampling strategy, summarized in Table 1 and Table 2. Another two models (conv2d6x22 and dws2d6x22) are trained with the same configuration except that the input to the first Conformer block is scaled by a factor of square root of the IO dimension described in (Vaswani et al. 2017).
@@
5.2 Performance
model w/ overflow model w/o overflow
multiplier multiplier
conv2d6x22 6.85% conv2d6 3.26%
dws2d6x22 6.85% dws2d6 0.25%
5.3 Energy
5.4 Numeric Stability
Figure 3: Distribution of the max value for chunk-based decoding between vanilla convolution and DWS in log scale.
Figure 4: Distribution of Layernorm’s input’s max value for chunk-based decoding in log scale.
model
WER (FP16)
WER (FP32)
conv2d6
4.45%
4.41%
dws2d6
4.55%
4.56%
conv2d6x22
4.57%
4.47%
dws2d6x22
4.57%
4.49%
conv2d6x22
4.76% 4.72%
+ modified Softmax
Table 4: WER comparison of FP16 and FP32
5.5 Quality
6 Conclusions
Through architectural, model-level, and numerical optimizations, we demonstrate that Conformer CTC ASR models are capable of running on resource-constrained devices such as mobile phones, and wearables. The optimizations preserve recognition accuracy while performing faster than real time and consuming lesser energy. Our theoretical findings of techniques in numerical stabilization is applicable to a wide range of deep learning models and computing tasks.

### v2

@@ Abstract
Conformer-Based Speech Recognition
On Extreme Edge-Computing Devices
Mingbin Xuthanks: Equal contribution.1, Alex Jinthanks: left Apple after paper submission.2, Sicheng Wang1, Mu Su1, Tim Ng1, Henry Mason1, Shiyi Han1, Zhihong Lei1, Yaqiao Deng1, Zhen Huang1, Mahesh Krishnamoorthy1 1Apple 2Centre for Quantum Technologies, NUS mingbinxu@apple.com, alex_jin@u.nus.edu, {sicheng_wang,mu_su,tim_ng,hmason,shan26,zlei,yaqiao_deng,zhen_huang,maheshk}@apple.com
@@ 1 Introduction
Conformer-based Gulati et al. (2020) end-to-end (E2E) automatic speech recognition (ASR) Yao et al. (2021); Zhang et al. (2022) with streaming capabilities He et al. (2019) have made numerous advances recently. This has paved the way for fully neural speech recognition on resource-constrained mobile devices. These systems also have numerous advantages over conventional hybrid-HMM ASR Hinton et al. (2012).
First, the training procedure is simplified; the entire system can be defined in a single deep learning framework such as PyTorch or TensorFlow. Second, recent work (e.g. Miao et al., 2019; Sainath et al., 2020; Li et al., 2020; Lei et al., 2023a, b) shows E2E ASR systems can provide better Word-Error-Rate (WER) when compared to conventional hybrid ASR systems. Third, with the continued advancement of deep learning applications, special hardware accelerators such as NVIDIA’s Graphics Processing Units (GPU), Google’s Tensor Processing Units (TPU), and Apple’s Neural Engine (ANE) are becoming increasingly popular. A fully neural ASR system can best utilize such hardware advancements and operate with high throughput while minimizing energy consumption.
In this paper, we present optimizations to enable fully E2E neural network based ASR system under resource-constrained environments, such as smartphones, wearables, and home automation devices. Operating fully offline saves cloud computing resources while providing stronger user privacy Xu et al. (2023) guarantees, as the user’s speech does not need to be transmitted outside of the device.
@@ 2 Prior Work
Improving the efficiency of the Transformer architecture has seen substantial interest. Tay et al. (2023) provides a comprehensive survey primarily concentrating on model architecture improvements. Kim et al. (2023) is another noteworthy resource which delves deeper into considerations specific to hardware configurations. Linear Transformer Katharopoulos et al. (2020) is a key technique, mitigating the computationally expensive softmax function Bridle (1989) within the attention mechanism. Softmax is also susceptible to numeric overflow problems when computing with limited numerical range. Hoffer et al. (2018); Zhang and Sennrich (2019) discuss alternative normalization methods other than Batchnorm Ioffe and Szegedy (2015) and Layernorm Ba et al. (2016) to improve computational efficiency and numerical stability in low precision environments. Principles for optimizing transformers have been described in Apple (2022) which target Apple hardware, but are generally applicable for similar devices. Within the domain of speech recognition, Squeezeformer Kim et al. (2022) stands as a seminal work focusing on efficiency optimization, particularly with respect to the Conformer architecture. The paper uses depthwise separable convolution subsampling to substantially save computation which is central to MobileNet Howard et al. (2017). It’s worth mentioning that the majority of prior work focuses on improving training efficiency by making modifications to the existing model architecture. As a result, these changes require model retraining to achieve efficiency improvements. In contrast, our research primarily concentrates on post-training, inference-only processes while avoiding model retraining whenever possible.
3 Backbone Model
Our backbone model is built upon the Conformer neural architecture Gulati et al. (2020) as shared acoustic encoder while connectionist temporal classification Graves et al. (2006) (CTC) and Attention-based Encoder Decoder (AED) Chan et al. (2016) as dual decoders trained with multitask learning mechanism Caruana (1997).
@@ 3 Backbone Model
Similar to prior work (e.g. Gulati et al., 2020), we stack transformer Vaswani et al. (2017) layers and convolution LeCun et al. (1998) layers alternatively to convert speech frames into high-level representation. We use a relative sinusoidal positional encoding Dai et al. (2019) into transformer layers. Since our goal is to stream ASR on edge devices, we adopt the chunk-based attention strategy to better balance accuracy and dependency of future audio frames Yao et al. (2021); Zhang et al. (2022).
4 Proposed Optimizations
@@ 4.1 Depthwise Separable Convolution
In the original Conformer encoder design Gulati et al. (2020), the subsampling module at the beginning of the architecture is implemented using two vanilla convolution layers. Our profiling shows that vanilla convolution subsampling accounts for 32.8% of the overall computation and becomes expensive on resource-constrained devices. To alleviate this bottleneck, we used the idea of depthwise separable convolution Howard et al. (2017); Chollet (2017) as a drop-in replacement and reduced this computational bottleneck to 4.0% whilst maintaining the WER Kim et al. (2022), making it particularly well-suited for inference tasks on mobile devices.
While most of the research emphasizes depthwise separable convolution’s (DWS) computational efficiency and small memory footprint, its effect on reducing dynamic range of the outputs needs more study. The possible reason could be that DWS reduces the number of multiply-accumulate operations needed for the convolution filters, hence the chance of bigger values. Low numeric range is of great importance for model deployment on edge devices equipped with hardware accelerators. Those hardware often operate in low precision (e.g.fp16) to ease the burden of storage and memory and are exposed to overflow.
@@ 4.2 Memory-aware Graph Execution
(a) Common compute flow of MHA
(b) ANE-optimized compute flow of MHA
Figure 1: $bz$, $h$ and $f$ refers to batch size, number of attention heads and feature dimension respectively, whereas $d=f/h$. Firstly, we transposed the input and output of Conformer CTC, expanding the input tensor to the desired shape of $(B,C,1,S)$. This transformation allowed us to execute most layers on the hardware accelerator as per Principle 1. Additionally, we extensively employed split and concatenation operations to enhance L2 cache residency (Principle 2). To address the issue of undesired memory copies resulting from batched matrix multiplication layers, we replaced them with Einstein summation operations (Principle 3).
In Apple’s white paper Apple (2022) on deploying transformers on the Apple Neural Engine (ANE), four principles are elaborated for optimizing transformers on the ANE:
@@
The key idea behind these 4 principles is being aware of high cost invoked by memory copies between CPU and our hardware accelerator. In our implementation, we adhered to the aforementioned principles. We demonstrate how to rewrite multihead attention (MHA) in Figure 1 as an example. More importantly, operations not supported by hardware accelerator were positioned at the beginning or end of the network graph, thus minimizing copies in the memory.
@@ 4.3 Stability of Layer Normalization
Layer normalization has become the de facto normalization method in transformers after Attention is all you need Vaswani et al. (2017). This normalization technique is widely used in the Conformer CTC architecture. On the other hand, modern hardware accelerators for deep learning often exploit lower precision compute paths in order to reduce memory and boost computation throughput. In the Conformer model, we observed that layer normalization and hardware accelerators are often in dissonance with each other. The reason is that skip connections in the Conformer model join values of varying magnitudes to a single tensor and this often leads to numerical underflows or overflows in low precision compute paths. For example, the maximum value is 65504 in half precision floating point format IEEE (2008). As a contrast, the maximum value is $383.4e$ 38 in single precision floating point format.
@@
$=\Big{(}\frac{|-\frac{S}{2}|}{\frac{1}{2}(\frac{2}{M})^{1/p}\sum_{i=1}^{n}|v_{i}|}\Big{)}^{p}+$ (3)
${\color[rgb]{1,1,1}hello}\Big{(}\frac{|\frac{S}{2}|}{\frac{1}{2}(\frac{2}{M})^{1/p}\sum_{i=1}^{n}|v_{i}|}\Big{)}^{p}$ (4)
@@
On a practical note, the pre-normalizer we used for our experiment was the one from Lemmas A1 and A2 (B) with $n=512$, which gave a slightly lower normalization constant than what Corollary 1 suggests. This worked well in our setup because attaining or even getting close to the maximum value as stated in Lemma 3 requires atypical distribution of values with very few extreme values and everything else being 0. This does not happen in practice, however, with the most common distribution of values observed being Gaussian.
@@ 4.4 Scaling of Softmax
Another common constraint on hardware accelerators is their limited support in complex operations. For example, hardware accelerators may choose to omit support for exponential operations Hu et al. (2018); Li et al. (2018). In such cases, we seek to implement such operations in memory instead, namely using lookup tables (LUT). However, since LUTs are slow and expensive in terms of memory consumption, we would like the tables to be as small as possible. To this end, we introduce a technique called conditional re-scaling for softmax layers:
@@
To interpret the above transformation, we first assume that our LUT gives reasonably accurate approximation for $x_{i}$’s below 4096. Next we take FP16 as an example of our low precision compute paths. We note that for values greater than 4096, gaps between values jump in increments of 4 according to IEEE 754-2008 IEEE (2008). Under such scenario, the softmax function behaves similarly to an argmax operation. Since gaps of values between 2048 and 4096 jump in increments of 2, the “argmax behavior" is largely preserved after the re-scaling and exponentiation.
@@
Figure 3: Energy consumption (in joules) for 200 queries of the original Conformer CTC vs Depthwise Separable Convolution (DWS) architectures. Blue and green bars represent the values on CPU and hardware accelerators, respectively. The y-axis is in log scale.
@@ 5.1 Setup
The training corpus contains 17k-hour audio-transcript pairs where the audio is randomly sampled from anonymized virtual assistant queries and human-annotated. We curate 20k queries in the same manner to form an accuracy test set. We use it to examine the accuracy of the optimizations. 200 queries are sampled from the accuracy test set and serve as the performance test set. The audio is decoded lightweightedly with CTC prefix beam search so as to rule out as many computationally intensive components as possible Graves et al. (2006). The data choice and the training recipe do not play important role in the experiments because the proposed methods focus on hardware acceleration. The experiments are conducted on iPhone XR and Apple Watch Series 7.
Two models (conv2d6 and dws2d6) are trained with the same hyper-parameters but minor difference in subsampling strategy, summarized in Appendix A. Another two models (conv2d6x22 and dws2d6x22) are trained with the same configuration except that the input to the first Conformer block is scaled by a factor of square root of the IO dimension described in Vaswani et al. (2017). Additionally we decode greedily on watch to show that encoder’s workload dominates.
5.2 Performance
model w/ overflow model w/o overflow
multiplier multiplier
conv2d6x22 6.85% conv2d6 3.26%
dws2d6x22 6.85% dws2d6 0.25%
5.3 Energy
5.4 Numeric Stability
Figure 4: Distribution of the max value between vanilla convolution and DWS in log scale.
Figure 5: Distribution of Layernorm’s input’s max value in log scale.
model WER (FP16) WER (FP32)
conv2d6 4.45% 4.41%
dws2d6 4.55% 4.56%
conv2d6x22 4.57% 4.47%
dws2d6x22 4.57% 4.49%
conv2d6x22 4.76% 4.72%
+ modified Softmax
Table 2: WER comparison of FP16 and FP32
5.5 Quality
6 Conclusions
Through architectural and numerical optimizations, we demonstrate that Conformer CTC ASR models are capable of running on resource-constrained devices such as mobile phones, and wearables. The optimizations preserve recognition accuracy while performing faster than real time and consuming lesser energy. Our theoretical findings of techniques in numerical stabilization is applicable to a wide range of deep learning models and computing tasks.
Appendix A Hyper Parameters
conv2d6x22
follows the recipe of Yao et al. (2021); Zhang et al. (2022), where the subsampling output is multiplied by $\sqrt{512}$ before being fed into the first conformer layer. The multiplier is originated from the transformer work Vaswani et al. (2017). Its hyper-parameters are summarized in Table 3.
dws2d6x22
is produced by replacing vanilla convolutional subsampling with depthwise separable convolution (DWS). Their difference is compared in Table 4.
conv2d6
is indentical to conv2dx22 except that multiplier is not applied.
dws2d6
is same as dws2dx22 but without applying the multiplier.
@@ Appendix A Hyper Parameters
Appendix B Mean Absolute Deviation Normalization on Example Distributions
Definition A1. A desirable low precision pre-normalizer maps a distribution of values to a bounded region, $[-C,C]$, for some small $C$.
Lemma A1. $f(\textbf{x})=\frac{\textbf{x}}{\frac{1}{n}\sum_{i=1}^{n}|x_{i}|}$ is a desirable low precision pre-normalizer for uniform distributions.
[…]
Thus, $f(\textbf{x})=\frac{2\textbf{x}}{L}\sim unif[-2,2]$.
Lemma A2. $f(\textbf{x})=\frac{\textbf{x}}{\frac{1}{n}\sum_{i=1}^{n}|x_{i}|}$ is a desirable low precision pre-normalizer for normal distributions.
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n}|x_{i}|$ $=\mathbb{E}[|\textbf{x}|]$
[…]
${\color[rgb]{1,1,1}zero}(by\ symmetry)$
$=\sqrt{\frac{2}{\pi}}\sigma.$